### v2

@@ 2.3 PCA and Sparse PCA (SPCA)
Principal Component Analysis (PCA) [47] is a well-known unsupervised feature extraction method. It aims to find an orthogonal linear transformation matrix that projects the original data vectors into a lower-dimensional manifold, maximizing the variance of the projected samples. The optimization problem of PCA can be formulated as the following self-contained regression form:
@@
To obtain sparse weights and improve interpretability, Sparse PCA (SPCA) [32] was proposed by adding ridge and lasso constraints to PCA:
@@
The self-contained regression term in SPCA incorporates an encoder-decoder architecture, selectively transforming a sparse subset of input features into encoded latent representations and reconstructing them through an orthogonal decoder, as Figure 2 shows. This approach gains flexibility by relaxing the transformation matrix $\mathbf{Q}$. When $\mathbf{Q}=\mathbf{U}$, SPCA reduces to PCA.
@@
By minimizing reconstruction errors, SPCA preserves the global data structure in an unsupervised manner, maintaining essential geometric relationships within the data manifold. The lasso constraint enables SPCA to assign weights only to the necessary features. The orthogonality constraint ensures that the latent features are complementary and reduces redundancy. Therefore, SPCA is well-suited for UFS tasks, particularly in balancing the clustering performance and feature interpretability.
@@ 2.4 Tensor-based unsupervised feature selection
Tensor-based unsupervised feature selection remains an emerging research area. Most of the existing methods rely on domain knowledge and are limited to certain application scenarios. For example, [20] proposes an image-transformation-based feature selection method, aiming at vehicle detection. [34] constructs a multi-omics data tensor and exploits feature selection criterion designed using the knowledge in bioinformatics. Recently, some methods designed for general cases have emerged. GRLTR [35] combines graph regularization with tensor tubal rank constraint to simultaneously preserve global and local information. However, it still requires vectorization. CPUFS [18] derives a model based on CPD and effectively preserves the tensor structure by minimizing the reconstruction error. It represents a promising approach to tensor-based UFS. However, CPUFS does not derive sparse factor matrices. Instead, it applies spectral clustering and utilizes a sparse linear classifier matrix to perform UFS, leading to insufficient utilization of the information in factor matrices and reduced interpretability of selected features. Additionally, the constructed graph in CPUFS needs the number of classes as prior knowledge and can be unreliable due to noise features.
Recently, developing sparse PCA models for UFS appears to be a new trend [36, 37, 38, 39]. PCA-like methods do not rely on the construction of graphs. By global manifold learning, they can preserve the data structure and achieve superior average performance. What’s more, they utilize the factor matrix that maximizes the variance of data samples in the projected space, whose sparse weights demonstrate a clearer relevance of the original features and have better interpretability. Unfortunately, existing PCA-like methods are non-tensor-based. Motivated by the above considerations, this study develops a sparse TPCA-based methodology tailored for UFS tasks.
@@ 3 Methodology
In this section, we extend the framework of SPCA to tensor form by utilizing techniques from T-SVDM and derive a model called Sparse Tensor PCA based on $\star_{\mathbf{M}}$ product (STPCA-MP). Then, we deduce a convex version of STPCA-MP for UFS tasks by merging the encoder and decoder tensors into a single sparse reconstruction tensor. We propose an efficient slice-by-slice optimization algorithm that individually solves each dual subproblem in the transform domain. Eventually, by considering both the transform matrix $\mathbf{M}$ and the learned reconstruction tensor, we compute a feature score map and perform feature selection. The overall framework is illustrated in Figure 1.
@@ 3.1 Model formulation of STPCA-MP
We formulate the model in the third-order tensor scenario for the simplicity of deduction and algorithm design. The principle can be extended to high-order scenarios in a recursive fashion. See Section 4 for details.
Given a data tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ containing $n$ samples of size $d_{1}\times d_{2}$, we can reorient it with an order set $D^{o}$ such that the second mode represents the sample dimension. We denote the rotated data tensor by $\mathcal{X}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$. Based on (8), the row dimensionality of each frontal slice in $\mathcal{X}_{D^{o}}$ can be reduced with $\mathcal{U}^{H}\in\mathbb{C}^{k\times q\times p}(k<d)$, followed by reconstruction with $\mathcal{U}$:
@@
Intuitively, we can first formulate a tensor PCA model following the idea of one-dimensional PCA in (9):
$\underset{\mathcal{U}}{\min}$ ${\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}$ (12)
$s.t.$ $\quad\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{U}=\mathcal{I}.$
Then, inspired by the encoder-decoder structure of SPCA, we use another tensor $\mathcal{Q}\in\mathbb{C}^{d\times k\times p}$ to perform feature extraction instead of $\mathcal{U}$. By doing so, we obtain a more flexible reconstruction framework that yields a better latent feature representation:
$\underset{\mathcal{U}}{\min}$ ${\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{Q}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}$ (13)
$s.t.$ $\quad\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{U}=\mathcal{I}.$
Specifically, $\mathcal{Q}$ contains the weights that measure the importance of each original feature during the projection into the latent feature space. To encode only the necessary original features, $\mathcal{Q}$ is constrained to have sparse loadings. Meanwhile, for highly relevant features, we hope the model can treat them equally, instead of preferring certain features, especially when $dm\gg n$. Therefore, lasso regression and ridge regression terms are introduced as constraints for $\mathcal{Q}$, resulting in a sparse Tensor PCA model:
$\underset{\mathcal{U},\mathcal{Q}}{\min}$ ${\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{Q}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}+\alpha\sum\limits^{pn}_{j=1}{\|\left(\mathbf{Q}_{(1)}\right)_{:j}\|^{2}_{2}}+\beta\sum\limits^{pn}_{j=1}{\|\left(\mathbf{Q}_{(1)}\right)_{:j}\|_{1}}$ (14)
$s.t.$ $\quad\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{U}=\mathcal{I}.$
By obtaining the optimal $\mathcal{Q}$, the sparse tensor principal components can be extracted from the data tensor.
Discussion: To this end, the proposed model is directly applicable to tasks traditionally addressed by SPCA (e.g., image compression, visualization, and denoising) while preserving the tensor structure. However, the encoder tensor alone is suboptimal for UFS tasks. In the following section, we reformulate the model to meet UFS-specific requirements.
@@ 3.2 Convex version of STPCA-MP for UFS
As previously discussed, the self-contained regression term has an encoder-decoder structure. Therefore, the importance of each original feature is not only expressed by the weights contained in $\mathcal{Q}$ (encoder) but also described by $\mathcal{U}$ (decoder). Comprehensively considering their relationship is essential for adapting the model to UFS.
Dual Problem: Since the transformation between tensor $\mathcal{A}$ and its counterpart $\hat{\mathcal{A}}$ is uniquely defined by $\hat{\mathcal{A}}=\mathcal{A}\times_{3}\mathbf{M}$, Problem (14) has its dual problem in the transform domain, which means they can be solved simultaneously. To facilitate optimization, we can rewrite Problem (14) in the transform domain where matrix algebra can be applied:
$\underset{\hat{\mathcal{U}},\hat{\mathcal{Q}}}{\min}$ $\sum\limits^{p}_{i=1}{\|\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}-\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}\|^{2}_{F}}+\lambda\sum\limits^{p}_{j=1}{\|\hat{\mathcal{Q}}_{::i}\|_{2,1}}$ (15)
$s.t.\quad\hat{\mathcal{U}}_{::i}^{H}\left(\hat{\mathcal{U}}_{::i}\right)=\mathbf{I},i\in[p].$
The $\ell_{2,1}$ norm is employed to simultaneously achieve the goals of lasso regression and ridge regression.
Variable Merging: Notably, there exists a tensor $\mathcal{A}\in\mathbb{C}^{d\times d}$ whose $i$-th frontal slice in the transform domain $\hat{\mathcal{A}}_{::i}=\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}$. Each row vector in $\hat{\mathcal{A}}_{::i}$ contains reconstruction weights that combine elements from both $\hat{\mathcal{U}}_{::i}$ and $\hat{\mathcal{Q}}^{H}_{::i}$. Therefore, $\mathcal{A}$ is a more suitable feature selection tensor than $\mathcal{Q}$. Guided by $D^{o}$, each frontal slice in $\hat{\mathcal{A}}$ reconstructs the fibers of a certain mode in the data tensor, utilizing the structural information hidden in the mode organization. Remarkably, it can be proven that the optimal $\hat{\mathcal{A}}$ is a Hermitian positive semidefinite (HPSD) matrix, detailed by Theorem 1.
@@
Norm Simplification: Since $\hat{\mathcal{A}}_{::i}=\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}$ and $d\gg k$, $\hat{\mathcal{A}}_{::i}$ is also a low-rank matrix. It is widely known that nuclear norm $\|\cdot\|_{*}$ is often used as the convex approximation to the low-rank matrix constraint. However, its high computational complexity poses a challenge. Fortunately, in our case, $\|\hat{\mathcal{A}}_{::i}\|_{*}={\rm Tr}\left(\hat{\mathcal{A}}_{::i}\right)$ due to the HPSD constraint, significantly reducing computational complexity.
Meanwhile, provided that $\hat{\mathcal{U}}_{::i}$ is orthogonal, it is easy to prove that
@@
Therefore we can use ${\|\hat{\mathcal{A}}_{::i}\|}_{2,1}$ to simplify optimization. Moreover, minimizing ${\|\hat{\mathcal{A}}_{::i}\|}_{2,1}$ and ${\rm Tr}\left(\hat{\mathcal{A}}_{::i}\right)$ is equivalent to minimizing ${\|\mathcal{A}_{::i}\|}_{2,1}$ and ${\rm Tr}\left(\mathcal{A}_{::i}\right)$ when $\mathbf{M}$ is orthogonal. This implies that the elements in a column vector of $\mathcal{A}_{::i}$ (representing different weights for the same feature) are minimized toward zero for redundant features, thereby fulfilling the feature selection task.
@@
Discussion: The convex version of STPCA-MP exhibits a compact model expression. By merging the sparse encoder and orthogonal decoder, the reconstruction tensor simultaneously assigns weights to original features and their latent representations, comprehensively measuring feature importance.

### v3

@@ 2.3 PCA and Sparse PCA (SPCA)
Principal Component Analysis (PCA) [50] is a fundamental unsupervised feature extraction method. It seeks an orthogonal linear transformation that projects high-dimensional data onto a lower-dimensional manifold while maximizing the variance of the projected samples. The PCA optimization problem can be expressed in the following self-contained regression form:
@@
To enhance interpretability through sparse feature weights, Sparse PCA (SPCA) [35] introduces ridge and lasso constraints to the PCA framework:
@@
The self-contained regression term in SPCA implements an encoder-decoder architecture:
  • Encoder ($\mathbf{Q}$): Selectively transforms a sparse subset of input features into latent representations
  • Decoder ($\mathbf{U}$): Reconstructs original data through an orthogonal transformation
as illustrated in Figure 2. This formulation enhances flexibility by decoupling the transformation matrix $\mathbf{Q}$ from $\mathbf{U}$. Notably, when $\mathbf{Q}=\mathbf{U}$, SPCA reduces to standard PCA.
@@
Through the minimization of reconstruction errors, SPCA preserves global data structures in an unsupervised manner while maintaining essential geometric relationships within the data manifold. The lasso constraint enables selective feature weighting by assigning non-zero weights only to necessary features. Simultaneously, the orthogonality constraint ensures complementarity of latent features and reduces redundancy. These properties make SPCA particularly well-suited for unsupervised feature selection (UFS), effectively balancing clustering performance and feature interpretability.
@@ 2.4 Tensor-based unsupervised feature selection
Tensor-based unsupervised feature selection (UFS) remains an emerging research domain where most existing methods rely heavily on domain-specific knowledge, limiting their applicability to particular scenarios. For instance, [23] proposes an image-transformation-based feature selection method specialized for vehicle detection, while [37] constructs multi-omics data tensors with feature selection criteria designed using bioinformatics knowledge. Although recent methods targeting general applications have emerged—such as GRLTR [38] which combines graph regularization with tensor tubal rank constraints to preserve global and local information—they still require vectorization. CPUFS [21] represents a promising tensor-based approach that preserves tensor structures via CP decomposition reconstruction, but it fails to derive sparse factor matrices. Instead, CPUFS applies spectral clustering and utilizes a sparse linear classifier matrix for feature selection, leading to insufficient utilization of factor matrix information and reduced interpretability. Furthermore, its graph construction requires prior knowledge of class numbers and remains vulnerable to noisy features.
Recently, sparse PCA models constitute a growing trend in UFS [39, 40, 41, 42] due to their ability to avoid graph construction dependencies, preserve data structures through global manifold learning, deliver superior average performance, and provide interpretable sparse weights via variance-maximizing factor matrices. However, existing PCA-based methods lack tensor formulations, limiting their applicability to higher-dimensional data. To bridge this gap, we develop a novel sparse TPCA-based methodology specifically designed for tensor UFS tasks.
@@ 3 Methodology
This section extends the SPCA framework to tensor formulation by integrating T-SVDM techniques, establishing a novel Sparse Tensor PCA based on $\star_{\mathbf{M}}$ product (STPCA-MP) model. We subsequently derive a convex variant of STPCA-MP for UFS applications by unifying the encoder and decoder tensors into a single sparse reconstruction tensor. To optimize this model efficiently, we develop a slice-by-slice algorithm that independently solves each dual subproblem in the transform domain. Finally, leveraging both the transformation matrix $\mathbf{M}$ and the learned reconstruction tensor, we compute feature score maps to perform feature selection. Figure 1 illustrates the complete framework, with key mathematical symbols summarized in Table 3.
Table 3: Symbols used throughout the deduction.
Notation Description
$\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ Data tensor containing $n$ samples
$\hat{\mathcal{X}}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ Data tensor in the transform domain
$\tilde{\mathcal{X}}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ Reconstructed data tensor
$D^{o}\in\mathbb{Z}^{3}$ Order set defining the order of modes
$\mathcal{X}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$ Data tensor permuted by $D^{o}$
$\mathcal{U}$, $\hat{\mathcal{U}}\in\mathbb{C}^{q\times k\times p}$ Decoder tensor for $\mathcal{X}_{D^{o}}$ in the original and the transform domain
$\mathcal{Q}$, $\hat{\mathcal{Q}}\in\mathbb{C}^{q\times k\times p}$ Encoder tensor for $\mathcal{X}_{D^{o}}$ in the original and the transform domain
$\mathcal{Q}_{::i}$ The $i$-th frontal slice of $\mathcal{Q}$
$\mathbf{Q}_{(1)}$ 1-mode unfolding of $\mathcal{Q}$
$\left(\mathbf{Q}_{(1)}\right)_{:j}$ The $j$-th column vector in $\mathbf{Q}_{(1)}$
$\mathcal{A}$, $\hat{\mathcal{A}}\in\mathbb{C}^{q\times q\times p}$ Reconstruction tensor for $\mathcal{X}_{D^{o}}$ in the original and the transform domain
$\mathcal{I}\in\mathbb{C}^{q\times q\times p}$ Identity tensor under $\star_{\mathbf{M}}$ product
$\alpha$ Regularization parameter of lasso regression
$\beta$ Regularization parameter of ridge regression
$\lambda$ Regularization parameter of $\ell_{2,1}$-norm
$\eta$ Regularization parameter of trace function
@@ 3.1 Model formulation of STPCA-MP
We formulate the model for third-order tensors to simplify deduction and algorithmic development, noting that the principle extends recursively to higher-order scenarios (see Section 4 for generalization details).
Consider a data tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ containing $n$ samples of dimensions $d_{1}\times d_{2}$. Through reorientation via an ordered set $D^{o}$ where the second mode corresponds to the sample dimension, we obtain the rotated tensor $\mathcal{X}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$. Note that following rotation, each frontal slice of $\mathcal{X}_{D^{o}}$ undergoes sample-wise mean centering along its second mode (sample dimension). Building on the T-SVDM framework (9), we reduce row dimensionality of each frontal slice in $\mathcal{X}_{D^{o}}$ using $\mathcal{U}^{H}\in\mathbb{C}^{k\times q\times p}$ ($k<d$), followed by reconstruction with $\mathcal{U}$:
@@
Intuitively, we can first formulate a tensor PCA model inheriting the principle of data reconstruction and orthogonality constraint in PCA (Problem (10)):
$\underset{\mathcal{U}}{\min}\quad{\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}$ (13)
[…]
which minimizes the data reconstruction error without compromising the tensor structure.
Building on the encoder-decoder architecture of SPCA [35], we introduce a distinct transformation tensor $\mathcal{Q}\in\mathbb{C}^{q\times k\times p}$ to replace $\mathcal{U}$ for feature extraction. This yields a more flexible reconstruction framework that enhances latent feature representation:
$\underset{\mathcal{U}}{\min}\quad{\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{Q}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}$ (14)
[…]
Similar to SPCA, Problem (14) adopts an encoder-decoder framework:
  • Tensor encoder ($\mathcal{Q}$): Projects input features into a latent space while preserving the intrinsic tensor structure and orientation relationships.
  • Tensor decoder ($\mathcal{U}$): Recovers the original tensor through a series of orthogonal transformations, ensuring minimal reconstruction error.
To enforce feature selection during encoding, we impose sparsity constraints on the transformation tensor $\mathcal{Q}$, ensuring only discriminative features are retained in the latent representation. For highly correlated features, particularly in high-dimensional scenarios ($d_{m}\gg n$), we impose equality constraints to prevent model bias toward specific features. This is achieved by incorporating lasso and ridge regression terms as $\mathcal{Q}$ constraints, yielding the sparse tensor PCA model:
$\underset{\mathcal{U},\mathcal{Q}}{\min}\quad{\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{Q}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}+\alpha\sum\limits^{pn}_{j=1}{\|\left(\mathbf{Q}_{(1)}\right)_{:j}\|^{2}_{2}}+\beta\sum\limits^{pn}_{j=1}{\|\left(\mathbf{Q}_{(1)}\right)_{:j}\|_{1}}$ (15)
[…]
which can be viewed as an extension of SPCA to a tensor form. Solving for the optimal $\mathcal{Q}$ enables extraction of sparse tensor principal components from the data tensor.
Discussion: To this end, the proposed model naturally extends SPCA applications to tensor-structured data, including traditional tasks such as image compression, visualization, and denoising, while preserving intrinsic structural relationships. However, the encoder tensor $\mathcal{Q}$ alone proves suboptimal for UFS due to insufficient feature discriminability. The following section reformulates the model specifically for UFS requirements.
@@ 3.2 Convex version of STPCA-MP for UFS
As previously established, the self-contained regression term embodies an encoder-decoder architecture. Consequently, feature importance is jointly encoded by $\mathcal{Q}$ (encoder) and $\mathcal{U}$ (decoder). Comprehensive integration of their relationship is essential for UFS adaptation.
Dual Problem: Given the bijective mapping $\hat{\mathcal{A}}=\mathcal{A}\times_{3}\mathbf{M}$ between tensors, Problem (15) admits a dual formulation in the transform domain. This duality enables simultaneous optimization. For computational efficiency, we recast the problem using matrix algebra:
$\underset{\hat{\mathcal{U}},\hat{\mathcal{Q}}}{\min}\quad\sum\limits^{p}_{i=1}{\|\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}-\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}\|^{2}_{F}}+\lambda\sum\limits^{p}_{j=1}{\|\hat{\mathcal{Q}}_{::i}\|_{2,1}}$ (16)
[…]
The $\ell_{2,1}$ norm concurrently satisfies lasso and ridge regression objectives. Critically, Problem (16) demonstrates tensor reconstruction via matrix outer products. Compared to vector outer products in one-dimensional SPCA, this formulation more effectively incorporates structural priors into sparse weight optimization.
Variable Merging: As Problem (16) is a non-convex problem requiring alternating update of $\hat{\mathcal{U}}$ and $\hat{\mathcal{Q}}$, a convex relaxation is beneficial for optimization. Consequently, we developed a convex formulation strategy through variable merging. As illustrated in Figure 3, tensor products in the transform domain are decomposed into multiple outer products of matrices. This enables pairwise multiplication of each frontal slice between the encoder and decoder, ultimately generating a novel reconstruction tensor. Each row vector in $\hat{\mathcal{A}}_{::i}$ integrates reconstruction weights from both $\hat{\mathcal{U}}_{::i}$ and $\hat{\mathcal{Q}}^{H}_{::i}$, establishing $\mathcal{A}$ as a superior feature selection tensor compared to $\mathcal{Q}$ alone. Guided by $D^{o}$, each frontal slice of $\hat{\mathcal{A}}$ reconstructs mode-specific fibers in the data tensor, thereby leveraging structural information inherent in the mode organization.
Theorem 1 demonstrates that the optimal $\hat{\mathcal{A}}$ constitutes a Hermitian positive semidefinite (HPSD) matrix.
@@
Figure 3: A schematic diagram of merging the encoder and decoder tensors.
Norm Simplification: Given $\hat{\mathcal{A}}_{::i}=\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}$ and $d\gg k$, each $\hat{\mathcal{A}}_{::i}$ is inherently low-rank. While the nuclear norm $\|\cdot\|_{*}$ serves as the standard convex approximation for low-rank constraints, its computational demands present significant challenges. Fortunately, the HPSD constraint enables the equivalence $\|\hat{\mathcal{A}}_{::i}\|_{*}={\rm Tr}\left(\hat{\mathcal{A}}_{::i}\right)$, dramatically reducing computational complexity.
Furthermore, leveraging the orthogonality of $\hat{\mathcal{U}}_{::i}$, we establish:
@@
This enables substitution with ${\|\hat{\mathcal{A}}_{::i}\|}_{2,1}$ to streamline optimization. Moreover, when $\mathbf{M}$ is orthogonal, minimizing ${\|\hat{\mathcal{A}}_{::i}\|}_{2,1}$ and ${\rm Tr}\left(\hat{\mathcal{A}}_{::i}\right)$ is equivalent to minimizing ${\|\mathcal{A}_{::i}\|}_{2,1}$ and ${\rm Tr}\left(\mathcal{A}_{::i}\right)$. Consequently, column vectors in $\mathcal{A}_{::i}$ (representing weights for individual features) are collectively minimized toward zero for redundant features, thereby fulfilling feature selection.
@@
The HPSD constraint, while emerging naturally from variable merging, plays a dual role: it statistically guarantees positive contributions from all projection directions to the reconstruction process. This property intrinsically suppresses random noise features, leading to enhanced weight interpretability and ultimately more effective feature selection.
Discussion: The convex STPCA-MP formulation achieves model compactness through variable merging. By unifying the sparse encoder and orthogonal decoder, the reconstruction tensor $\mathcal{A}$ simultaneously weights original features and their latent representations, providing a comprehensive feature importance measure.